Past attendance records show that the probability that the chairman of the board attends a meeting is 0.65, that the president of the company attends a meeting is 0.9, and that they both attend a meeting is 0.6. Would you say that the chairman and the president act independently
regarding their attendance at the meetings of the board of directors? Why?
1
Expert's answer
2019-10-04T10:32:54-0400

The events A and B are independent if and only if (iff)


"P(A\cap B)=P(A)\cdot P(B)"


Let "A" be the event, that the chairman of the board attends a meeting, "B" be the event, that the president of the company attends a meeting.

Given that "P(A)=0.65, P(B)=0.9, P(A\cap B)=0.6"

Check the independence


"P(A)\cdot P(B)=0.65\cdot0.9=0.585\not=0.6=P(A\cap B)"

Therefore, the chairman and the president don't act independently regarding their attendance at the meetings of the board of directors.